two coins are tossed let T the number of tails that occurs determine the values of the random variable T

1
Expert's answer
2022-03-31T02:25:55-0400

Let "h" denote a head and "t" denote a tail. Then, possible events for 2 coins are: "ht", "th", "hh", "tt". The random variable "T", which is defined on space "\Omega=\{ht,th,hh,tt\}" , takes on values: "T(ht)=1, T(th)=1, T(hh)=0,T(tt)=2." The probabilities are as follows: "P(T=2)=P(tt)=0.25", "P(T=1)=P(th,ht)=0.5", "P(T=0)=P(hh)=0.25."